# Aligning Cyber Space with Physical World: A Comprehensive Survey on Embodied AI

### v7

@@ VI-B Embodied Task Planning
Traditional embodied task planning methods are usually based on explicit rules and logical reasoning. For example, symbolic planning algorithms such as STRIPS [310] and PDDL[311], and search algorithms like MCTS [312] and A* [313], are used to generate plans. However, these methods often rely on predefined rules, constraints, and heuristics that are rigid and may not adapt well to dynamic or unforeseen changes in the environment. With the popularity of LLMs, many works have attempted to use LLMs for planning or to combine traditional methods with LLMs, leveraging the rich world knowledge embedded within them for reasoning and planning without the need for handcrafted definitions, greatly enhancing the model’s generalization capabilities.
VI-B1 Planning utilizing the Emergent Capabilities of LLMs
@@
Before the scale-up of natural language models, task planners were similarly implemented by training models like BERT on embodied instruction datasets such as Alfred [314] and Alfworld [315], as demonstrated by FILM [316]. However, this approach was limited by the examples in the training set and could not effectively align with the physical world. Nowadays, thanks to the emergent capabilities of LLMs, they can decompose abstract tasks using their internal world knowledge and chain-of-thought reasoning, similar to how humans reason through task completion steps before acting. For example, Translated LM [317] and Inner Monologue [318] can break down complex tasks into manageable steps and devise solutions using their internal logic and knowledge systems without additional training, such as ReAct [319]. Similarly, the multi-agent collaboration framework ReAd [320] is proposed for efficient self-refinement of plans via the different prompts. Additionally, some approaches abstract past successful examples into a series of skills stored in a memory bank to consider during inference and improve planning success rates [321, 322, 323]. And some works utilize code as the reasoning medium instead of natural language. where task planning is generated as code based on the available API library[324, 325, 326]. Furthermore, multi-turn reasoning can effectively correct potential hallucinations in task planning, a focus of many LLM-based agent studies. For instance, Socratic Models [327] and Socratic Planner [328] use Socratic questioning to derive reliable planning.
@@
VI-B2 Planning utilizing the visual information from embodied perception model
Based on the above discussion, it is particularly important to further integrate visual information into task planning (or replanning). In this process, object labels, locations, or descriptions provided by visual input can offer critical references for task decomposition and execution by LLMs. Through visual information, LLMs can more accurately identify target objects and obstacles in the current environment, thereby optimizing task steps or modifying subtask objectives. Some works use an object detector to query the objects present in the environment during task execution and feed this information back to the LLM, allowing it to modify unreasonable steps in the current plan [329, 327, 330]. RoboGPT considers the different names of similar objects within the same task, further improving the feasibility of replanning [10]. However, the information provided by labels is still too limited. Can further scene information be provided? SayPlan [331] proposes using hierarchical 3D scene graphs to represent the environment, effectively mitigating the challenges of task planning in large, multi-floor, and multi-room settings. Similarly, ConceptGraphs [332] also adopts 3D scene graphs to provide environmental information to LLMs. Compared to SayPlan, it offers more detailed open-world object detection and presents task planning in a code-based format, which is more efficient and better suited to the demands of complex tasks.
@@ VI-B2 Planning utilizing the visual information from embodied perception model
VI-B3 Planning utilizing the VLMs
@@
However, task planning is only the first step for an agent in completing an instruction task; subsequent action planning determines whether the task can be accomplished. In the experiments from RoboGPT [10], the accuracy of task planning reached 96%, but the overall task completion rate was only 60%, limited by the performance of the low-level planner. Therefore, whether an embodied agent can transition from the cyber space of “imagining how tasks are completed” to the physical world of “interacting with the environment and completing tasks” hinges on effective action planning.
VI-C Embodied Action Planning
Section VI-B discusses the definitions and differences between task planning and action planning. It is evident that action planning must address real-world uncertainties because the granularity of subtasks provided by task planning is insufficient to guide agents in environmental interaction. Generally, agents can achieve action planning in two ways: 1) using pre-trained embodied perception and embodied intervention models as tools, incrementally completing the subtasks specified by task planning through APIs, 2) utilizing the VLA model’s inherent capabilities to derive action planning. Furthermore, the results of the action planner’s execution are fed back to the task planner to adjust and improve task planning.
VI-C1 Action utilizing APIs
A typical approach involves providing LLMs with the definitions and descriptions of various well-trained policy models as context, enabling them to understand these tools and determine how and when to invoke them for specific tasks [299, 329]. Additionally, by generating code, a series of more granular tools can be abstracted into a function library for invocation rather than directly passing the parameters needed for sub-tasks to navigation and grasping models [326]. Given the uncertainty of the environment, Reflexion can further adjust these tools during execution to achieve better generalization [336]. Optimizing these tools can enhance the robustness of the agent, and new tools may be required to complete unknown tasks. DEPS, under the premise of zero-shot learning, endows LLMs with various role settings to learn diverse skills while interacting with the environment. During subsequent interactions, LLMs can learn to select and combine these skills to develop new ones [337].
This hierarchical planning paradigm allows agents to focus on high-level task planning and decision-making while delegating specific action execution to policy models, simplifying the development process. The modularity of the task planner and action planner enables independent development, testing, and optimization, enhancing system flexibility and maintainability. This method allows the agent to adapt to various tasks and environments by invoking different action planners and facilitating modifications without requiring significant changes to the agent’s structure. However, invoking external policy models may introduce latency, potentially impacting response time and efficiency, especially in real-time tasks. The agent’s performance is critically dependent on the quality of the policy models. If the policy models are ineffective, the overall performance of the agent will suffer.
VI-C2 Action utilizing VLA model
@@ VII-A Embodied World Model
Sim-to-Real involves creating world models in simulation that closely resemble real-world environments, helping algorithms generalize better when transferred. The world model approach aims to build an end-to-end model that maps vision to action, or even anything to anything, by predicting the next state in a generative or predictive manner to make decisions. The biggest difference between such world models and VLA models is that VLA models are first trained on large-scale internet datasets to achieve high-level emergent capabilities and then co-finetuned with real world robot data. In contrast, world models are trained from scratch on physical world data, gradually developing high-level capabilities as the amount of data increases. However, they remain low-level physical world models, somewhat akin to the mechanism of human neural reflex systems. This makes them more suitable for scenarios where both inputs and outputs are relatively structured, such as autonomous driving (input: vision, output: throttle, brake, steering wheel) or object sorting (input: vision, instructions, numerical sensors, output: grasping the target object and placing it in the target location). They are less suited for generalization to unstructured, complex embodied tasks.
Figure 15: Embodied world models can be roughly divided into three type. (a) Generation-based Methods train to learn the transformation relationship between the input space and the output space using an autoencoder framework. (b) Prediction-based Methods can be seen as a more general framework where a world model is trained in latent space. (c) Knowledge-driven Methods inject artificially constructed knowledge into the model, giving the model world knowledge to obtain output that meets the given knowledge constraints. Note that the components within the dashed line are optional.
TABLE XI: Summary of the embodied world methods discussed in VII-A.
Type Method Years Main tasks
Generation-based World Models[338] 2018 Car Racing
Sora [17] 2024 Video Generation
Pandora[339] 2024 Real-time Controllable Video Generation
3D-VLA[340] 2024 Embodied Reasoning and Localization, Multimodal Goal Generation, Robot Planning
DWM[341] 2024 D4RL Offline RL
Prediction-based I-JEPA[16] 2023 Visual Representation Learning, Image Classification, Object Counting, Depth Prediction
MC-JEPA[342] 2023 Visual Representation Learning, Optical Flow Estimation, Instance Segmentation, Video Segmentation
A-JEPA[343] 2023 Audio Representation Learning, Audio and Speech Classification
IWM[344] 2024 Visual Representation Learning, Image Classification, Image Segmentation
iVideoGPT[345] 2024 Video Prediction, Visual Planning, Visual Model-based RL
IRASim[346] 2024 Trajectory to Video Generation
STP[347] 2024 Robotic Motor Control
MuDreamer[348] 2024 DeepMind Visual Control Suite, Natural Background Setting
Knowledge-driven Lionel et al. [349] 2023 Probabilistic Reasoning, Relational Reasoning, Perceptual and Physical Reasoning, Social Reasoning
ElastoGen[350] 2024 4D Elastodynamics Generation
Liu et al.[351] 2024 Single-image 3D Reconstruction
Holodeck[71] 2024 3D Environments Generation
LEGENT[352] 2024 3D Environments Generation
GRUtopia [353] 2024 Language Instruction, Task Generation, Navigation, Social Interaction, Manipulation
Learning world models is promising of the physical simulation field. Compared to traditional simulation methods, it offers significant advantages, such as the ability to reason about interactions with incomplete information, meet real-time computation requirements, and improve prediction accuracy over time. The predictive capability of such world models is crucial, enabling robots to develop the physical intuition necessary to operate in the human world. As shown in Fig. 15, according to the learning pipeline of the world environment, they can be divided into Generation-based methods, Prediction-based methods and Knowledge-driven methods. We briefly summarize the methods mentioned in Table XI.
@@ VII-A1 Generation-based Methods
As the scale of models and data progressively increases, generative models have demonstrated the ability to understand and generate images (e.g., World Models [338]), videos (e.g., Sora [17], Pandora [339]), point clouds (e.g., 3D-VLA [340]) or other formats of data (e.g., DWM [341]) that conform to physical laws. This ability indicates that generative models can learn and internalize world knowledge. Specifically, after being exposed to vast amounts of data, generative models can not only capture the statistical properties of the data but also simulate the physical and causal relations of the real world through their intrinsic structures and mechanisms. Therefore, these generative models can be considered more than simple pattern recognition tools: they exhibit characteristics of world models. Consequently, the world knowledge embedded in generative models can be leveraged to enhance the performance of other models. By mining and utilizing the world knowledge represented in generative models, we can improve model generalization and robustness. This approach not only enhances the model’s adaptability to new environments but also increases its predictive accuracy on unknown data [340, 339]. However, generative models also have certain limitations and drawbacks. For instance, when data distribution is significantly biased or training data is insufficient, generative models may produce inaccurate or distorted outputs. Additionally, the training process for these models typically requires substantial computational resources and time, and the models often lack interpretability, which complicates their practical application. Overall, while generative models have shown great potential in understanding and generating content that conforms to physical laws, several technical and practical challenges must be addressed for their effective application. These challenges include improving model efficiency, enhancing interpretability, and addressing issues related to data bias. With ongoing research and development, generative models are expected to demonstrate even greater value and potential in future applications.
@@ VII-A2 Prediction-based Methods
The prediction-based world model predicts and understands the environment by constructing and utilizing internal representations. By reconstructing corresponding features in the latent space based on provided conditions, it captures deeper semantics and associated world knowledge. This model maps input information to a latent space and operates within that space to extract and utilize high-level semantic information, thereby enabling the robots to perceive the essential representation of the world environment (e.g., I-JEPA[16], MC-JEPA[342], A-JEPA[343], Point-JEPA [354], IWM[344]) and more accurately perform embodied downstream tasks (e.g., iVideoGPT[345], IRASim[346]), STP[347], MuDreamer[348]). Compared to pixel-level information, latent feature can abstract and decouple various forms of knowledge, allowing the model to handle complex tasks and scenes more effectively and improve its generalization capability[355]. For instance, in spatiotemporal modeling, the world model needs to predict the post-interaction state of an object based on its current state and the nature of the interaction, combining this information with its internal knowledge. Specifically, an embodied world model generates dynamic predictions of the environment by integrating perceptual information and prior knowledge. This approach relies not only on sensory data but also on inherent world knowledge to infer and predict environmental changes, thereby producing more accurate spatiotemporal predictions[345, 347, 348]. This process considers both the current state of objects and their historical data and contextual information.
Figure 16: The illustration of demonstration data collection. The yellow box on the left features operational demonstrations of the Franka and WidowX robotic arms, while human demonstrations are shown in the blue box. In the middle, the yellow box showcases operational scenarios of the UR5e and Franka robotic arms in simulation environments, and the blue box displays labeled simulation data. The purple box on the right presents the data formats of these datasets.
@@ VII-B Data Collection and Training
For sim-to-real adaptation, the high-quality data is important. Traditional data collection methods involve expensive equipment, precise operations, and are time-consuming and labor-intensive, often lacking flexibility. Recently, some efficient and cost-effective methods have been proposed for high-quality demonstration data collection and training. This section will discuss various methods for data collection in both real-world and simulated environments. Fig. 16 presents demonstration data from both real-world and simulated environments.
@@ VII-B1 Real-World Data
Training large, high-capacity models on high-volume, rich datasets has demonstrated remarkable capabilities and significant success in effectively addressing downstream applications. For instance, LLMs such as ChatGPT, GPT-4, and LLaMA have not only excelled in the field of NLP but have also provided excellent problem-solving capabilities for downstream tasks. Therefore, is it possible to train an embodied large model in the robotics field, which possesses strong generalization capabilities through training and can adapt to new scenarios and robotic tasks. This requires a large volume of embodied datasets to provide data for model training. Open X-Embodiment [303] is an embodied dataset from 22 different robots, with 527 skills and 160,266 tasks. The collected data consists of authentic demonstration data from robots, obtained by recording the process of executing operations. This dataset primarily focuses on domestic and kitchen settings, involving items such as furniture, food, and tableware. The operations are mainly centered around pick-and-place tasks, with a small portion involving more complex maneuvers. The high-capacity model RT-X trained on this dataset demonstrates excellent transfer capabilities. UMI [358] proposed a data collection and policy learning framework. They designed a handheld gripper and an elegant interface for data collection, enabling portable, low-cost, and information-rich data collection for challenging bimanual and dynamic demonstration data. By simply modifying the training data, the robot can achieve zero-shot generalizable, bimanual, and precise tasks. Mobile ALOHA [359] is a low-cost full-body mobile manipulation system. It can be used to collect task data for bimanual operations under full-body mobility, such as frying shrimp and serving dishes. Training agents with data collected by this system and static ALOHA can improve the performance of mobile manipulation tasks. Such agents can serve as home assistants or work assistants. In human-agent collaboration [360], humans and agents learn together during data collection, reducing human workload, accelerating data acquisition, and improving data quality. Specifically, in an embodied scenario, during data collection, humans provide initial action inputs. Subsequently, the agent refines these actions through iterative perturbation and denoising processes, gradually optimizing them to produce precise, high-quality operational demonstrations. This entire process can be summarized as follows: humans contribute intuition and diversity in operations, while agents handle optimization and stability, reducing reliance on operators, enabling execution of more complex tasks, and gathering higher-quality data.
@@ VII-B2 Simulated Data
The aforementioned data collection methods involve directly collecting demonstration data in the real world for agent training. Such collection methods often require significant manpower, material resources, and time, leading to inefficiency. Therefore, in most cases, researchers can choose to collect datasets in simulation environments for model training. Collecting data in simulation environments does not require extensive resources and can generally be automated by programs, saving a lot of time. CLIPORT [294] and Transporter Networks [361] collected demonstration data from the Pybullet simulator for end-to-end network model training and successfully transferred the models from simulation to real world. GAPartNet [362] constructed a large-scale part-centric interactive dataset GAPartNet, providing rich part-level annotations for perception and interaction tasks. They proposed a pipeline for domain-generalized 3D part segmentation and pose estimation, which can generalize well to unseen object categories in both simulators and the real world. SemGrasp [289] built a large-scale grasping text-aligned dataset CapGrasp, which is a semantically rich dexterous hand grasping dataset from virtual environments.

### v8

@@ VI-A Embodied Task Planning
Traditional embodied task planning methods are usually based on explicit rules and logical reasoning. For example, symbolic planning algorithms such as PDDL[173], and search algorithms like MCTS [174] and A* [175], are used to generate plans. However, these methods often rely on predefined rules, constraints, and heuristics that are rigid and may not adapt well to dynamic or unforeseen changes. With the popularity of LLMs, many works have attempted to use LLMs for planning or to combine traditional methods with LLMs, leveraging the rich embedded world knowledge for reasoning and planning without the need for handcrafted definitions, greatly enhancing the model’s generalization capabilities.
VI-A1 Planning utilizing the Emergent Capabilities of LLMs
@@
Before the scale-up of natural language models, task planners were similarly implemented by training models like BERT on embodied instruction datasets such as Alfred [176] and Alfworld [177], as demonstrated by FILM [178]. However, this approach was limited by the examples in the training set and could not effectively align with the physical world. Nowadays, thanks to the emergent capabilities of LLMs, they can decompose abstract tasks using their internal world knowledge and chain-of-thought reasoning, similar to how humans reason through task completion steps before acting. For example, Translated LM [179] and Inner Monologue [180] can break down complex tasks into manageable steps and devise solutions using their internal logic and knowledge systems without additional training. Similarly, the multi-agent collaboration framework ReAd [181] efficiently self-refined plans via different prompts. Additionally, some approaches abstract past successful examples into a series of skills stored in a memory bank to consider during inference and improve planning success rates [182, 183, 184]. Some works utilized code as the reasoning medium instead of natural language, where task planning is generated as code based on the available API library[185, 186]. Furthermore, multi-turn reasoning can effectively correct potential hallucinations in task planning. For instance, Socratic Models [187] and Socratic Planner [188] used Socratic questioning to derive reliable planning.
@@
VI-A2 Planning utilizing the visual information from embodied perception model
Based on the above discussion, it is important to integrate visual information into task planning (or replanning). In this process, object labels, locations, or descriptions provided by visual input can offer critical references for task decomposition and execution by LLMs. Through visual information, LLMs can more accurately identify target objects and obstacles in the current environment, thereby optimizing task steps or modifying subtask objectives. Some works use an object detector to query the objects present in the environment during task execution and feed this information back to the LLM, allowing it to modify unreasonable steps in the current plan [189, 187, 190]. RoboGPT considers the different names of similar objects within the same task, further improving the feasibility of replanning [8]. However, the information provided by labels is still too limited. Can further scene information be provided? SayPlan [191] proposes using hierarchical 3D scene graphs to represent the environment, effectively mitigating the challenges of task planning in large, multi-floor, and multi-room settings. Similarly, ConceptGraphs [192] also adopts 3D scene graphs to provide environmental information to LLMs. Compared to SayPlan, it offers more detailed open-world object detection and presents task planning in a code-based format, which is more efficient and better suited to the demands of complex tasks.
@@ VI-A2 Planning utilizing the visual information from embodied perception model
VI-A3 Planning utilizing the VLMs
@@
However, task planning is only the first step for an agent in completing an instruction task. Subsequent action planning determines whether the task can be accomplished. In the experiments from RoboGPT [8], the accuracy of task planning reached 96%, but the overall task completion rate was only 60%, limited by the performance of the low-level planner. Therefore, whether an embodied agent can transition from the cyber space of “imagining how tasks are completed” to the physical world of “interacting with the environment and completing tasks” hinges on effective action planning.
VI-B Embodied Action Planning
The distinction between task planning and action planning highlights that action planning must address real-world uncertainties due to the insufficient granularity of task planning subtasks [198]. Action planning can be achieved by: 1) using pre-trained embodied models to complete subtasks via APIs, or 2) leveraging the VLA model’s capabilities. The results from action planning are fed back to refine task planning.
VI-B1 Action utilizing APIs
A common approach involves providing LLMs with definitions of well-trained policy models to understand and use them effectively for specific tasks [199, 189]. By generating code, LLMs can abstract tools into a function library, allowing better handling of sub-tasks [186]. Reflexion adjusts these tools during execution to improve generalization [200]. DEPS enables LLMs to learn and combine various skills through zero-shot learning [201]. The hierarchical planning paradigm simplifies development by separating high-level task planning from specific action execution through policy models. This modularity allows independent development, testing, and optimization, enhancing flexibility and maintainability. While it enables adaptability to various tasks and environments, reliance on external policy models can introduce latency and affect performance, making the quality of these models crucial for overall agent effectiveness.
VI-B2 Action utilizing VLA model
@@
VI-B3 Scalability in Diverse Environments
Scalability in embodied agents involves adapting to increased complexity in larger and more diverse environments through robust perception, efficient decision-making, and resource optimization. Strategies include hierarchical SLAM for mapping, multi-modal perception, and energy-efficient edge computing. Collaborative scalability is enhanced via multi-agent systems and decentralized communication, while generalization relies on domain adaptation to operate in new environments.
@@ VII-A Embodied World Model
Sim-to-Real involves creating simulation-based world models that closely resemble real-world environments. These models predict the next state to make decisions and are trained from scratch on physical world data, unlike VLA models which are pre-trained on large-scale datasets and fine-tuned with real-world data. World models are effective for structured tasks like autonomous driving and object sorting but are less suited for unstructured, complex tasks.
Learning world models is promising of the physical simulation field. Compared to traditional simulation methods, it offers significant advantages, such as the ability to reason about interactions with incomplete information, meet real-time computation requirements, and improve prediction accuracy over time. The predictive capability of such world models is crucial, enabling robots to develop the physical intuition necessary to operate in the human world. As shown in Fig. 10, according to the learning pipeline of the world environment, they can be divided into Generation-based methods, Prediction-based methods and Knowledge-driven methods.
@@ VII-A1 Generation-based Methods
As the scale of models and data increases, generative models have demonstrated the ability to understand and generate images (e.g., World Models [203]), videos (e.g., Sora [16], Pandora [204]), point clouds (e.g., 3D-VLA [205]) or other formats of data (e.g., DWM [206]) that conform to physical laws. This capability suggests that generative models can internalize world knowledge. Specifically, after exposure to large datasets, these models not only capture statistical properties but also simulate physical and causal relationships through their intrinsic structures. Thus, generative models function as more than just pattern recognition tools; they exhibit characteristics of world models. The embedded world knowledge in these models can be harnessed to enhance the performance of other models. By leveraging this knowledge, we can improve model generalization, robustness, adaptability to new environments, and predictive accuracy on unseen data [205, 204]. However, generative models also have certain limitations and drawbacks. For instance, they may produce inaccurate or distorted outputs when faced with biased data distributions or insufficient training data. Moreover, these models require substantial computational resources and time for training and often lack interpretability, which hinders practical application. While generative models have shown promise in generating content that adheres to physical laws, challenges such as improving efficiency, enhancing interpretability, and mitigating data bias must be addressed for broader application. Continued research is likely to unlock further value and potential in these models.
Figure 10: Embodied world models can be roughly divided into three type. (a) Generation-based Methods learn the transformation relation between the input space and the output space using an autoencoder framework. (b) Prediction-based Methods are more general frameworks where a world model is trained in latent space. (c) Knowledge-driven Methods inject artificially constructed knowledge into the model, giving the model world knowledge to obtain output that meets the given knowledge constraints. Note that the components within the dashed line are optional.
@@ VII-A2 Prediction-based Methods
The prediction-based world model predicts and understands the environment by constructing and utilizing internal representations. By reconstructing corresponding features in the latent space based on provided conditions, it captures deeper semantics and associated world knowledge. This model maps input information to a latent space and operates within that space to extract and utilize high-level semantic information, thereby enabling the robots to perceive the essential representation of the world environment (e.g., I-JEPA[15], MC-JEPA[207], A-JEPA[208], Point-JEPA [209], IWM[210]) and more accurately perform embodied downstream tasks (e.g., iVideoGPT[211], IRASim[212]), STP[213], MuDreamer[214]). Latent features, unlike pixel-level information, can abstract and decouple various forms of knowledge, enabling models to handle complex tasks and scenes more effectively while enhancing generalization [215]. In spatiotemporal modeling, for instance, a world model predicts an object’s post-interaction state by integrating its current state, the nature of the interaction, and internal knowledge. Specifically, embodied world models generate dynamic environmental predictions by combining perceptual information with prior knowledge, relying on both sensory data and inherent world knowledge to accurately infer and predict environmental changes [211, 213, 214]. This process considers the current state of objects alongside their historical and contextual information.
@@ VII-A3 Knowledge-driven Methods
VII-A4 Limitations
Current limitations of world models include handling the complexity and variability of real-world environments, such as high-dimensional sensory inputs, dynamic and stochastic elements, and long-term dependencies. Many models struggle with generalization across unseen scenarios, often requiring extensive training data and computational resources. Additionally, sim-to-real transfer remains problematic, as simulated environments fail to fully capture real-world physics and noise. These limitations can be addressed by integrating more realistic simulators, incorporating multimodal sensory inputs, and using hierarchical or modular architectures. Improving data efficiency, enhancing transfer learning techniques, and incorporating real-world priors can also enable more accurate predictions and adaptive decision-making. Recently, the Cosmos Platform [223] was proposed by NVIDIA, which contains autoregressive and diffusion models for Text-to-World and Video-to-World generation, to accelerate the development of physical AI systems. It may give us some inspirations about how to build an embodied world model.
@@ VII-B Data Collection and Training
For sim-to-real adaptation, the high-quality data is important. Traditional data collection methods involve expensive equipment and precise operations, which are time-consuming and labor-intensive. Recently, some efficient and cost-effective methods have been proposed for high-quality demonstration data collection and training. We discuss data collection methods in both real-world and simulated environments.
@@ VII-B1 Real-World Data
Training large, high-capacity models with rich datasets has shown great success. This approach is also promising for robotics, where large, diverse datasets can enhance generalization and adaptability. Open X-Embodiment [202] provides data from 22 robots with 527 skills and 160,266 tasks in domestic settings. UMI [224] offers a framework for collecting dynamic, bimanual data using a handheld gripper. Mobile ALOHA [225] enables data collection for full-body mobile manipulation tasks. Human-agent collaboration [226] improves data quality and efficiency by combining human input with agent refinement processes.
@@ VII-B2 Simulated Data
Data collection in real-world settings is resource-intensive and time-consuming, making simulation-based data collection an attractive alternative. Simulated environments allow for automated, efficient data collection. For example, CLIPORT [168] and Transporter Networks [227] utilized Pybullet simulator data for training and successfully transferred models to real-world applications. GAPartNet [228] developed a large-scale dataset with detailed part-level annotations for better object interaction in both simulations and reality. SemGrasp [167] created the CapGrasp dataset for semantically rich hand grasping in virtual environments.